Assume the quotes indicated below are from a dealer in the NY currency market. Forward prices are quoted as the absolute forward premium or discount (not percentage) in basis points (one basis point = .0001).

Part 1. Forward exchange rates

1 a. If the dealer’s spot market quotes for the Canadian dollar are 1.3218 1.3222, and the dealer’s 9-month forward quotes for the Canadian dollar are: -81 -77, what are the dealer’s effective 9-month forward bid and ask prices for the Canadian dollar?

b. If a customer were to enter a 9-month forward contract with this dealer to sell $1m to obtain Canadian dollars, how many Canadian dollars would the customer receive in nine months?

2 a. If the dealer’s spot market quotes for the Russian ruble are: 65.7120 65.7130, and the dealer’s 3-month forward quotes for the ruble are: 2003 2040, what are the dealer’s effective 3-month forward bid and ask prices for the Russian ruble?

b. If a customer were to enter a 3-month forward contract with this dealer to sell ₽1m to obtain US dollars, how many US dollars would the customer receive in three months?

3. If forward prices are used as an indicator of the likely increase or decrease in value of a currency, which of the foreign currencies above are expected to increase in value over the indicated time period, and which are expected to decrease in value?
Answer #1
1.a) The outright quotes for the CD = 1.3218-.0081;1.3222-.0077
giving the rates of 1.3137/1.3145
The bid price is 1.3137 and the ask price is 1.3145
1.b) The ask price should be used. The CDs that would be
received are 1000000*1.3145 = 1314500 CD
2.a) Effective bid rate = 65.7120+0.2003 = $                       65.9123
Effective ask rate = 65.7130+0.2040 = $                       65.9170
2-b) The bid price should be used. The amount of $ to be received = 1000000/65.9123 = $                   15,171.68
3) CD is expected to depreciate and $ is expected to appreciate for [1] above.
$ is expected to depreciate and Rouble is expected to depreciate for [2] above.